### v1

@@ 2.5 Mapping to activity patterns in the human brain
We further explored the biological plausibility of LLM-derived conceptual representations by mapping them onto the activity patterns in the human brain. Using fMRI data from the THINGS-data collection [43], we fitted a voxel-wise linear encoding model to predict neural responses evoked by viewing concept images, based on the corresponding conceptual representations derived from LLaMA3-70B (Methods 4.9). Fig. 7a–c shows the prediction performance maps, indicating voxels where the predicted activations best correlated with actual activations ($P<0.01$, FDR controlled). The LLM-derived conceptual representations successfully predicted activity patterns across widely distributed brain regions, encompassing the visual cortex and beyond. In particular, category-selective regions were more strongly represented, including the lateral occipital complex (LOC), occipital face area (OFA), fusiform face area (FFA), parahippocampal place area (PPA), extrastriate body area (EBA) and medial place area (MPA). These patterns are consistent with prior work suggesting that abstract semantic information was primarily represented in the higher-level visual cortex [53, 54, 55, 56]. Meanwhile, significant prediction performance was observed in early visual regions, including V1, indicating that information processed in low-level visual areas is also relevant and can be effectively inferred by LLMs trained exclusively on language data. The consistent location of informative voxels across participants supports the generality of our findings.
@@
As shown in Fig. 7d–f, LLM-derived conceptual representations and the similarity embedding shared a considerable amount of explained variance, particularly within higher-level visual areas, with some overlap in early visual regions. This indicates that both accounted well for neural responses associated with visual concepts. However, the analysis of unique variance revealed that LLM-derived conceptual representations captured a greater proportion of variance, extending from higher-level regions to early visual areas including V1 (Fig. 7d). By comparison, the unique variance explained by the similarity embedding was primarily concentrated in low-level regions, with limited contributions in higher-level areas such as hV4 and EBA (Fig. 7e). This difference was even more pronounced when leveraging the full dimensionality of LLM-derived conceptual representations for analysis (Supplementary Information SI 1.6). These findings suggest that the brain’s encoding of concepts transcends mere similarities, with some information better captured by the conceptual representations formed by LLMs in context. Nevertheless, certain aspects of visual information relevant to human behaviour may not be adequately represented in these representations learned solely from language.
@@
Figure 7: Prediction performance of LLM-derived conceptual representation (LLaMA3-70B) and comparison with baseline models in voxel-wise encoding. ac, Prediction performance of LLM-derived conceptual representation visualized on cortical maps for three individual participants. Only voxels with statistically significant correlations between predicted and actual activations are coloured (FDR $P<0.01$). df, Comparison between LLM-derived conceptual representation and a similarity embedding learned from human similarity judgments. d, Variance uniquely explained by the LLM-derived conceptual representation. e, Variance uniquely explained by the similarity embedding. f, Shared variance explained by both models. gi, Comparison between LLM-derived conceptual representation and a static word embedding (fastText). g, Variance uniquely explained by the LLM-derived conceptual representation. h, Variance uniquely explained by the static word embedding. i, Shared variance explained by both models. The colours represent the proportion of explained variance, normalized relative to the noise ceiling and averaged over five runs.
@@ 3 Discussion
How are concepts represented and organized in the mind? Recent developments in LLMs [20, 57] provide a novel approach to this question. Here, we demonstrated that next-token prediction over language naturally gives rise to human-like conceptual representations and organization, even without real-world grounding. Our work builds upon the long-explored idea of vectors as conceptual representations [46, 7, 16, 42], while previous work has predominantly focused on word embeddings [52, 30]. We viewed concepts as latent representations used for word generation and guided LLMs to infer them from definitional descriptions. Our findings revealed that LLMs can adaptively derive concepts based on contextual demonstrations, reflecting the interrelationships among them. These representations converged towards a context-independent relational structure predictive of the models’ concept inference capacity, suggesting that language prediction inherently fosters the development of a shared conceptual structure. This structure supports the generalization of knowledge by effectively capturing key properties of human concepts, such as similarity judgments, categorical distinctions and gradient scales along various features. Notably, these conceptual representations showed a strong alignment with neural coding patterns observed in the human brain, even in response to non-linguistic stimuli like visual images. These findings suggest that LLMs serve as valuable tools for modelling human conceptual organization. By providing insights into the computational mechanisms underlying conceptual representation, this work establishes a foundation for enhancing the alignment between AI systems and human cognition.
@@
Concepts are considered as mental representations that abstract away specific details and enable flexible generalization in novel situations [3, 17, 4]. Using the reverse dictionary task, we showed that LLMs can effectively derive such conceptual representations from definitional descriptions. While traditional word embeddings have shown potential to capture certain properties required for conceptual representations [58, 52, 39, 59], their capacity is constrained by the inherent context-sensitivity of words, which do not correspond to concepts in a straightforward way [3, 60]. Accordingly, word embeddings are either limited by their static nature—failing to account for context-dependent nuances—or their contextual variability, which precludes consistent mapping to distinct concepts. In contrast, the conceptual representations derived from informative descriptions bypass the ambiguity of words and can be consistently mapped to appropriate terms despite contextual variations (Fig. 2). We argue that these representations are abstract, as their relational structures were consistently preserved across varying contexts (Fig. 3a–b). This consistency indicates that the representations capture the underlying relationships among concepts while discarding surface-level details, a hallmark of abstraction. Such abstraction is essential for generalization, enabling the learned relationships to be flexibly adapted to novel situations. Similar abstract, though disentangled, representations have been observed in humans [61], monkeys [62], rodents [63], and neural networks trained for multitasking [64]. However, the focus on task-specific low-level features is insufficient to model the richness of broadly generalizable real-world concepts. Our results, spanning diverse LLM architectures (Fig. 3c–d), reveal that abstract representations encompassing a wide array of real-world concepts can emerge solely from language prediction. This highlights a promising pathway towards modelling the complexity of the human conceptual repertoire.
@@
Relationships among concepts have long been a cornerstone of cognitive theories [46, 7, 14]. Here, we showed that language prediction naturally gives rise to interrelated concepts. This is demonstrated through the concept inference behaviour of LLMs, which was shaped by contextual cues rather than occurring in isolation (Fig. 2), and the consistently preserved relational structure within their representation spaces (Fig. 3a–b). Comparisons with human data further revealed that these relationships aligned with psychological measures of similarity and encapsulated a wealth of human conceptual knowledge (Fig. 4 and Fig. 5). According to conceptual role semantics (CRS) [65, 66], the meaning of a concept is determined by its relationships to other concepts and its role in thinking, rather than reference to the real world. It has been claimed that LLMs may possess human-like meaning in this sense, with language serving as a valuable source for inferring how people use concepts in thought [41]. However, evidence was needed to determine whether the objective of language prediction can lead to the discovery of the right conceptual roles. Our results support this claim, showing that the relationships among LLM-derived conceptual representations approximate human meaning. While these relationships could be further refined to match those of humans, particularly with respect to real-world grounding, compositionality and abstract reasoning [30, 67], our findings suggest that LLMs offer a prospective computational footing for implementing conceptual role semantics, paving the way for the development of more human-like, concept-based AI systems.
@@
Theories of concepts have identified various properties that conceptual representations must satisfy, highlighted by the symbolism-connectionism divide. Our results show that LLM-derived conceptual representations successfully reconcile the seemingly competing properties, integrating the definitions, relations and structures emphasized by the symbolic approaches with the graded and continuous nature of neural networks. These representations were structured in a way that their relationships supported straightforward computations of human-like similarities, categories (Fig. 4, Supplementary Information SI 1.4) and gradient distinctions (Fig. 5). Previous work has employed distributed word embeddings to approximate these aspects of human concept usage, revealing preliminary correspondence [50, 58, 52]. However, these embeddings have exhibited inconsistent alignment with human similarity judgments across various datasets [30]. Our results align with previous findings, showing that word embeddings primarily reflect association or relatedness but fail to capture genuine similarity (Fig. 4a–b, Supplementary Information SI 1.3). In contrast, LLM-derived conceptual representations demonstrated distinctive strengths. They excelled in modelling human similarity judgments, including those based on images, suggesting that they advantageously capture conceptual information independent of stimulus modality. These representations also supported flexible, context-dependent reorganization to reflect gradient scales along different features, surpassing word embeddings in both breadth and depth. Importantly, they exhibited superior alignment with neural activity patterns in the human brain (Fig. 7). Collectively, these findings underscore the potential of LLM-derived conceptual representations to advance our understanding of human concepts and to bridge the long-standing divide between symbolic and connectionist approaches.
@@
Vector-based models have been argued to plausibly capture the neural activations underlying actual cognitive processes in the brain [16, 8]. The mapping between LLM-derived conceptual representations and brain activity patterns suggests that they provide a plausible basis for analysing neural representations of concepts (Fig. 7). Compared to word embeddings and representations derived from human similarity judgments, the LLM-derived representations exhibited notable advantages, especially in accounting for neural activity patterns in high-level visual regions (Fig. 7d–i) that are associated with abstract semantic information [53, 54, 56]. However, they fell short in capturing certain aspects of behaviourally relevant, low-level visual information (Fig. 7d–f). To understand this discrepancy, we regressed the similarity representations onto the LLM-derived representations and found the most pronounced gaps in factors related to colour, followed by texture and shape, which are primarily perceptual properties (Supplementary Information SI 1.6, Fig. S11). This result aligns with previous research showing that blind individuals primarily acquire such properties through inference and differ significantly from sighted people in their knowledge of colour [40, 70]. Such information, therefore, may be inefficient to learn from language or even absent altogether. Despite these limitations, the substantial alignment between LLM-derived representations and neural responses to visual stimuli adds to the evidence that language prediction can orient models towards a shared representational structure of reality, even in the absence of real-world grounding [39, 68]. While prior work has hinted at this convergence [71], our findings explicitly characterize the conceptual structure emerging solely from language, setting the stage for future research on the alignment and interaction between linguistic and visual systems.
@@ 4.2 Details on deriving conceptual representations
We used data from the THINGS database [45] to probe LLMs’ conceptual representations via the reverse dictionary task. The dataset includes 1,854 concrete and nameable object concepts, paired with their WordNet synset IDs, definitional descriptions, several linked images and category membership labelled by humans. The concepts and images were selected to be representative of everyday objects commonly used in American English, providing a useful testbed for analysing model representations. To assess the generality of our results, we extended our analysis to a broader range of descriptions and concepts. Specifically, (1) we tested LLMs’ generalizability across different word classes (nouns, verbs, adjectives and adverbs), age-of-acquisition, and degrees of concreteness using a broader set of 21,402 concepts. The concepts were selected from the intersection of the three datasets: age-of-acquisition [97], concreteness [98] and Wordnet [99]. (2) We evaluated the consistency of LLMs’ predictions using three additional distinct definitional descriptions for each of the 1,854 THINGS concepts, which were generated by GPT-3.5 and manually checked for diversity (see Supplementary Information SI 2.1 for details). (3) We examined LLMs’ sensitivity to linguistic structure by introducing varying degrees of word order permutations to the query descriptions. We shuffled $30\%$, $60\%$ and $100\%$ of the words in the query descriptions from the THINGS database and reinserted them into the original text.
@@ 4.3 Structural analysis of representation spaces
An appropriate similarity function is needed for the (dis)similarity matrix to characterize the relational structure of the representation space. While cosine similarity has been widely adopted since the advent of static distributed word embeddings, it may be suboptimal for capturing non-linear relationships and sensitive to outliers [103]. We thus employed two different metrics including the cosine similarity and Spearman’s rank correlation. Comparison with human behavioural data suggests that the cosine similarity captures semantic similarity (e.g., cups and mugs) while Spearman’s rank correlation reflects both similarity and association (e.g., cups and coffee) (Supplementary Information SI 1.3, Fig. S3a–b). We primarily used Spearman’s rank correlation as the similarity function, with results from alternative metrics provided in Supplementary Information SI 1.2.
@@ 4.6 Categorization
For the categorization experiment, we employed high-level human-labelled natural categories in THINGS [45]. We removed subcategories of other categories, concepts belonging to multiple categories and categories with fewer than ten concepts [51]. This resulted in 18 out of 27 categories, including animal, body part, clothing, container, electronic device, food, furniture, home decor, medical equipment, musical instrument, office supply, part of car, plant, sports equipment, tool, toy, vehicle and weapon. These categories comprise 1,112 concepts.
@@ 4.8 Word embeddings for comparison
To validate the effectiveness of LLM-derived conceptual representations in capturing human knowledge and to explore their distinct advantages over traditional static word embeddings, we compared them with a state-of-the-art 300-dimensional static word embedding trained through fastText on Common Crawl and Wikipedia [49]. Unlike LLM representations, we used cosine similarity as the similarity measure for the word embeddings, as it consistently aligned better with human behavioural data (Supplementary Information SI 1.3).
The word embeddings were compared with human behavioural data using the same method applied to LLM conceptual representations across all experiments, except for the analysis of gradient distinctions along various features. As static word embeddings do not support context-dependent computations, we followed the procedure in previous work [52] to compute a scale vector based on several antonym pairs denoting opposite values of the target feature. For instance, the opposite values for the feature “size” were represented by $\left(\overrightarrow{\textrm{large}},\overrightarrow{\textrm{big}},\overrightarrow{\textrm{huge}}\right)$ and $\left(\overrightarrow{\textrm{small}},\overrightarrow{\textrm{little}},\overrightarrow{\textrm{tiny}}\right)$, with the scale vector $\overrightarrow{\prime\prime\textrm{size}\prime\prime}$ calculated as the average of the $3\times 3=9$ pairwise vector differences between the antonyms. The word embedding for each item was then projected onto this scale vector, and the resulting projections were correlated with human ratings for evaluation.

### v3

@@ 2.5 Mapping to activity patterns in the human brain
We further explored the biological plausibility of LLM-derived conceptual representations by mapping them onto the activity patterns in the human brain. Using fMRI data from the THINGS-data collection [40], we fitted a voxel-wise linear encoding model to predict neural responses evoked by viewing concept images, based on the corresponding conceptual representations derived from LLaMA3-70B (Methods 4.9). Fig. 7a–c shows the prediction performance maps, indicating voxels where the predicted activations best correlated with actual activations ($P<0.01$, FDR controlled). The LLM-derived conceptual representations successfully predicted activity patterns across widely distributed brain regions, encompassing the visual cortex and beyond. In particular, category-selective regions were more strongly represented, including the lateral occipital complex (LOC), occipital face area (OFA), fusiform face area (FFA), parahippocampal place area (PPA), extrastriate body area (EBA) and medial place area (MPA). These patterns are consistent with prior work suggesting that abstract semantic information was primarily represented in the higher-level visual cortex [50, 51, 52, 53]. Meanwhile, significant prediction performance was observed in early visual regions, including V1, indicating that information processed in low-level visual areas is also relevant and can be effectively inferred by LLMs trained exclusively on language data. The consistent location of informative voxels across participants supports the generality of our findings (Fig. S13a–c).
@@
As shown in Fig. 7d–f, LLM-derived conceptual representations and the similarity embedding shared a considerable amount of explained variance, particularly within higher-level visual areas, with some overlap in early visual regions. This indicates that both accounted well for neural responses associated with visual concepts. However, the analysis of unique variance revealed that LLM-derived conceptual representations captured a greater proportion of variance, extending from higher-level regions to early visual areas including V1 (Fig. 7d). By comparison, the unique variance explained by the similarity embedding was primarily concentrated in low-level regions, with limited contributions in higher-level areas such as hV4 and EBA (Fig. 7e). This difference was even more pronounced when leveraging the full dimensionality of LLM-derived conceptual representations for analysis (Supplementary Information SI 1.6). These findings suggest that the brain’s encoding of concepts transcends mere similarities, with some information better captured by the conceptual representations formed by LLMs in context. Nevertheless, certain aspects of visual information relevant to human behavior may not be adequately represented in these representations learned solely from language.
@@
Figure 7: Prediction performance of LLM-derived conceptual representation (LLaMA3-70B) and comparison with baseline models in voxel-wise encoding. ac, Prediction performance of LLM-derived conceptual representation visualized on cortical maps for three individual participants. Only voxels with statistically significant correlations between predicted and actual activations are colored (FDR $P<0.01$). df, Comparison between LLM-derived conceptual representation and a similarity embedding learned from human similarity judgments. d, Variance uniquely explained by the LLM-derived conceptual representation. e, Variance uniquely explained by the similarity embedding. f, Shared variance explained by both models. gi, Comparison between LLM-derived conceptual representation and a static word embedding (fastText). g, Variance uniquely explained by the LLM-derived conceptual representation. h, Variance uniquely explained by the static word embedding. i, Shared variance explained by both models. The colors represent the proportion of explained variance, normalized relative to the noise ceiling and averaged over five runs.
@@ 3 Discussion
In this paper, we demonstrated that next-token prediction over language naturally gives rise to human-like conceptual representations and organization, even without real-world grounding. Our work builds upon the long-explored idea of vectors as conceptual representations [43, 7, 16, 39], while previous work has predominantly focused on word embeddings [49, 27]. We viewed concepts as latent representations used for word generation and guided LLMs to infer them from definitional descriptions. Our findings revealed that LLMs can adaptively derive concepts based on contextual demonstrations, reflecting the interrelationships among them. These representations converged towards a context-independent relational structure predictive of the models’ concept inference capacity, suggesting that language prediction inherently fosters the development of a shared conceptual structure. This structure supports the generalization of knowledge by effectively capturing key properties of human concepts, such as similarity judgments, categorical distinctions and gradient scales along various features. Notably, these conceptual representations showed a strong alignment with neural coding patterns observed in the human brain, even in response to non-linguistic stimuli like visual images. These findings suggest that LLMs serve as promising tools for understanding human conceptual organization. By providing insights into the computational mechanisms underlying conceptual representation, this work establishes a foundation for enhancing the alignment between AI systems and human cognition.
@@
Concepts are considered as mental representations that abstract away specific details and enable flexible generalization in novel situations [3, 17, 4]. Using the reverse dictionary task, we showed that LLMs can effectively derive such conceptual representations from definitional descriptions. While traditional word embeddings have shown potential to capture certain properties required for conceptual representations [54, 49, 36, 55], their capacity is constrained by the inherent context-sensitivity of words, which do not correspond to concepts in a straightforward way [3, 56]. Accordingly, word embeddings are either limited by their static nature—failing to account for context-dependent nuances—or their contextual variability, which precludes consistent mapping to distinct concepts. In contrast, the conceptual representations derived from informative descriptions bypass the ambiguity of words and can be consistently mapped to appropriate terms despite contextual variations (Fig. 2). We argue that these representations are abstract, as their relational structures were consistently preserved across varying contexts (Fig. 3a–b). This consistency indicates that the representations capture the underlying relationships among concepts while discarding surface-level details, a hallmark of abstraction. Such abstraction is essential for generalization, enabling the learned relationships to be flexibly adapted to novel situations. Similar abstract, though disentangled, representations have been observed in humans [57], monkeys [58], rodents [59], and neural networks trained for multitasking [60]. However, the focus on task-specific low-level features is insufficient to model the richness of broadly generalizable real-world concepts. Our results, spanning diverse LLM architectures (Fig. 3c–d), reveal that abstract representations encompassing a wide array of real-world concepts can emerge solely from language prediction. This highlights a promising pathway towards modeling the complexity of the human conceptual repertoire.
@@
Relationships among concepts have long been a cornerstone of cognitive theories [43, 7, 14]. Here, we showed that language prediction naturally gives rise to interrelated concepts. This is demonstrated through the concept inference behavior of LLMs, which was shaped by contextual cues rather than occurring in isolation (Fig. 2), and the consistently preserved relational structure within their representation spaces (Fig. 3a–b). Comparisons with human data further revealed that these relationships aligned with psychological measures of similarity and encapsulated a wealth of human conceptual knowledge (Fig. 4 and Fig. 5). According to conceptual role semantics (CRS) [61, 62], the meaning of a concept is determined by its relationships to other concepts and its role in thinking, rather than reference to the real world. It has been claimed that LLMs may possess human-like meaning in this sense, with language serving as a valuable source for inferring how people use concepts in thought [38]. However, evidence was needed to determine whether the objective of language prediction can lead to the discovery of the right conceptual roles. Our results support this claim, showing that the relationships among LLM-derived conceptual representations approximate human meaning. While these relationships could be further refined to match those of humans, particularly with respect to real-world grounding, compositionality and abstract reasoning [27, 63], our findings suggest that LLMs offer a prospective computational footing for implementing conceptual role semantics, paving the way for the development of more human-like, concept-based AI systems.
@@
Theories of concepts have identified various properties that conceptual representations must satisfy, highlighted by the symbolism-connectionism divide. Our results show that LLM-derived conceptual representations successfully reconcile the seemingly competing properties, integrating the definitions, relations and structures emphasized by the symbolic approaches with the graded and continuous nature of neural networks. These representations were structured in a way that their relationships supported straightforward computations of human-like similarities, categories (Fig. 4, Supplementary Information SI 1.4) and gradient distinctions (Fig. 5). Previous work has employed distributed word embeddings to approximate these aspects of human concept usage, revealing preliminary correspondence [47, 54, 49]. However, these embeddings have exhibited inconsistent alignment with human similarity judgments across various datasets [27]. Our results align with previous findings, showing that word embeddings primarily reflect association or relatedness but fail to capture genuine similarity (Fig. 4a–b, Supplementary Information SI 1.3). In contrast, LLM-derived conceptual representations demonstrated distinctive strengths. They excelled in modeling human similarity judgments, including those based on images, suggesting that they advantageously capture conceptual information independent of stimulus modality. These representations also supported flexible, context-dependent reorganization to reflect gradient scales along different features, surpassing word embeddings in both breadth and depth. Importantly, they exhibited superior alignment with neural activity patterns in the human brain (Fig. 7). Collectively, these findings underscore the potential of LLM-derived conceptual representations to advance our understanding of human concepts and to bridge the long-standing divide between symbolic and connectionist approaches.
@@
Vector-based models have been argued to plausibly capture the neural activations underlying actual cognitive processes in the brain [16, 8]. The mapping between LLM-derived conceptual representations and brain activity patterns suggests that they provide a plausible basis for analyzing neural representations of concepts (Fig. 7). Compared to word embeddings and representations derived from human similarity judgments, the LLM-derived representations exhibited notable advantages, especially in accounting for neural activity patterns in high-level visual regions (Fig. 7d–i) that are associated with abstract semantic information [50, 51, 53]. However, they fell short in capturing certain aspects of behaviorally relevant, low-level visual information (Fig. 7d–f). To understand this discrepancy, we regressed the similarity representations onto the LLM-derived representations and found the most pronounced gaps in factors related to color, followed by texture and shape, which are primarily perceptual properties (Supplementary Information SI 1.6, Fig. S18). This result aligns with previous research showing that blind individuals primarily acquire such properties through inference and differ significantly from sighted people in their knowledge of color [37, 66]. Such information, therefore, may be inefficient to learn from language or even absent altogether. Despite these limitations, the substantial alignment between LLM-derived representations and neural responses to visual stimuli adds to the evidence that language prediction can orient models towards a shared representational structure of reality, even in the absence of real-world grounding [36, 64]. While prior work has hinted at this convergence [67], our findings explicitly characterize the conceptual structure emerging solely from language, setting the stage for future research on the alignment and interaction between linguistic and visual systems.
@@ 4.2 Details on deriving conceptual representations
We used data from the THINGS database [42] to probe LLMs’ conceptual representations via the reverse dictionary task. The dataset includes 1,854 concrete and nameable object concepts, paired with their WordNet synset IDs, definitional descriptions, several linked images and category membership labeled by humans. The concepts and images were selected to be representative of everyday objects commonly used in American English, providing a useful testbed for analyzing model representations. To assess the generality of our results, we extended our analysis to a broader range of descriptions and concepts. Specifically, (1) we tested LLMs’ generalizability across different word classes (nouns, verbs, adjectives and adverbs), age-of-acquisition, and degrees of concreteness using a broader set of 21,402 concepts. The concepts were selected from the intersection of the three datasets: age-of-acquisition [93], concreteness [94] and Wordnet [95]. (2) We evaluated the consistency of LLMs’ predictions using three additional distinct definitional descriptions for each of the 1,854 THINGS concepts, which were generated by GPT-3.5 and manually checked for diversity (see Supplementary Information SI 2.1 for details). (3) We examined LLMs’ sensitivity to linguistic structure by introducing varying degrees of word order permutations to the query descriptions. We shuffled $30\%$, $60\%$ and $100\%$ of the words in the query descriptions from the THINGS database and reinserted them into the original text.
@@ 4.3 Structural analysis of representation spaces
An appropriate similarity function is needed for the (dis)similarity matrix to characterize the relational structure of the representation space. While cosine similarity has been widely adopted since the advent of static distributed word embeddings, it may be suboptimal for capturing non-linear relationships and sensitive to outliers [99]. We thus employed two different metrics including the cosine similarity and Spearman’s rank correlation. Comparison with human behavioral data suggests that the cosine similarity captures semantic similarity (e.g., cups and mugs) while Spearman’s rank correlation reflects both similarity and association (e.g., cups and coffee) (Supplementary Information SI 1.3, Fig. S10a–b). We primarily used Spearman’s rank correlation as the similarity function, with results from alternative metrics provided in Supplementary Information SI 1.2.
@@ 4.6 Categorization
For the categorization experiment, we employed high-level human-labeled natural categories in THINGS [42]. We removed subcategories of other categories, concepts belonging to multiple categories and categories with fewer than ten concepts [48]. This resulted in 18 out of 27 categories, including animal, body part, clothing, container, electronic device, food, furniture, home decor, medical equipment, musical instrument, office supply, part of car, plant, sports equipment, tool, toy, vehicle and weapon. These categories comprise 1,112 concepts.
@@ 4.8 Word embeddings for comparison
To validate the effectiveness of LLM-derived conceptual representations in capturing human knowledge and to explore their distinct advantages over traditional static word embeddings, we compared them with a state-of-the-art 300-dimensional static word embedding trained through fastText on Common Crawl and Wikipedia [46]. Unlike LLM representations, we used cosine similarity as the similarity measure for the word embeddings, as it consistently aligned better with human behavioral data (Supplementary Information SI 1.3).
The word embeddings were compared with human behavioral data using the same method applied to LLM conceptual representations across all experiments, except for the analysis of gradient distinctions along various features. As static word embeddings do not support context-dependent computations, we followed the procedure in previous work [49] to compute a scale vector based on several antonym pairs denoting opposite values of the target feature. For instance, the opposite values for the feature “size” were represented by $\left(\overrightarrow{\textrm{large}},\overrightarrow{\textrm{big}},\overrightarrow{\textrm{huge}}\right)$ and $\left(\overrightarrow{\textrm{small}},\overrightarrow{\textrm{little}},\overrightarrow{\textrm{tiny}}\right)$, with the scale vector $\overrightarrow{\prime\prime\textrm{size}\prime\prime}$ calculated as the average of the $3\times 3=9$ pairwise vector differences between the antonyms. The word embedding for each item was then projected onto this scale vector, and the resulting projections were correlated with human ratings for evaluation.